# On the observation of cosmic strings via gravitational-wave lensing

### v2

@@ Abstract
We present a framework for detecting gravitational-wave signals lensed by cosmic strings (CSs), addressing a key gap in current searches. CSs, whose detection would provide a unique probe of high-energy physics and the early Universe, possess distinct topological and geometric features that require a dedicated search strategy. Our approach employs a full-wave transmission factor, expressed analytically via Fresnel integrals, which captures the characteristic diffraction and interference effects of the conical spacetime around a straight CS. We contrast CS lensing with the well-studied point mass lens (PML) model, highlighting their fundamental differences: CS lensing depends on cosmological distances, string tension $\Delta$, and wavelength $\lambda$, and produces two non-amplified images set by the global conical geometry. In contrast, PML lensing is governed by the distance-independent ratio $\sim M_{Lz}/\lambda$, where $M_{Lz}$ represents the redshifted mass of the lens, with image properties derived from the lens equation. For BBH mergers lensed by CSs, we show that the waveforms exhibit a characteristic beating pattern or time-separated, exact replicas. We derive a detectability bound on the string tension and, using Bayesian model selection, demonstrate that CS lensing is distinguishable from both unlensed and PML-lensed signals across a wide region of parameter space.
@@ I Introduction
Cosmic strings (CSs) are one-dimensional topological defects predicted to arise during spontaneous symmetry-breaking phase transitions in the early Universe, within a wide range of quantum field theories, including grand unified theories and string theory [1, 2, 3, 4, 5]. Their detection would provide a unique observational window into high-energy physics and the early Universe, potentially validating fundamental theoretical models. Advanced LIGO [6], Advanced Virgo [7], and KAGRA [8] observatories form an international network of ground-based gravitational-wave (GW) detectors. Together, they constitute the LIGO–Virgo–KAGRA (LVK) Collaboration, which has recently undertaken its fourth observing run (O4) [9]. The continuous improvement of detector sensitivity and global coordination among these observatories significantly enhance the prospects of detecting a broad range of GW sources, including potential signals from CSs. The LVK Collaboration has been actively searching for GW signals emitted by CS loops and kinks [10, 11]. When a network of CSs is assumed, their collective emission may contribute to the stochastic GW background [12, 13, 14, 15, 16]. In addition, several studies have explored the possibility of identifying individual GW bursts from CSs [17, 18].
@@
To address this gap, we introduce a comprehensive framework tailored for detecting gravitational lensing signatures from CSs. Central to our approach is a full-wave transmission factor that incorporates interference and diffraction effects from the conical spacetime geometry of CSs, providing a compact analytical expression in terms of Fresnel integrals readily available in scientific libraries such as Python’s scipy. This formulation enables the efficient construction of matched-filter templates and their straightforward implementation in data-analysis pipelines, laying the groundwork for targeted searches of CS-induced lensing effects in GW data. Focusing on the case of BBH mergers lensed by straight CSs, we characterize the resulting observational signatures, assess their detectability with current and future GW observatories, and examine potential mismodelling biases, matched-filtering strategies, and consistency tests to provide a quantitative evaluation of their impact on lensing detectability. Furthermore, we study the characteristic imprint left by CS lensing on detected GW and its discernibility from other, more common GW sources by computing the odds ratio PML lensed waveforms or unlensed waveforms.
This paper is organized as follows. In Sec. II, we review the main aspects of wave-optics effects in gravitational lensing, distinguishing between compact lenses and cosmic strings, and introduce the key parameters that characterize each regime. In Sec. III, we begin with the conical space-time geometry and present the full-wave solution for the transmission factor, parametrized by two dimensionless quantities: the scaled GW frequency—set by the string tension—and the source offset relative to the line of sight. Unlike compact lenses, the thin-lens approximation does not apply to cosmic strings; no lens equation is required, as image positions are dictated directly by the global conical geometry and the string tension. Finally, we compare cosmic string lensing with the point mass lens, emphasizing the fundamentally different physical mechanisms and observational signatures. Sec. IV discusses the observational imprints of cosmic string lensing and their detectability. In Sec. V, we analyze potential selection biases, including matched filtering strategies and consistency checks, and provide a quantitative assessment of their impact on lensing detectability. Finally, Sec. VI summarizes our conclusions. Additional technical details and estimate of the detection rate are provided in the appendices.
@@ III The transmission factor
III.1 Conical space-time
Suppose that the string (referred to as the “lens”) is positioned very close to the line of sight between the source of GW and the detector. For simplicity, we assume that the distance $d_{LS}$ between the GW source and the string is much longer than $d_{L}$ – the distance between the string and the detector. In this case the incident wave is just a plane wave and the reduced distance $\mathcal{D}\approx d_{L}$. Additionally, the string is considered to be orthogonal to the line of sight. This assumption does not preclude the application of our analysis to the possibility of inclination.
Consider the spacetime metric for a static cylindrically symmetric cosmic string lying along the $z$-axis [60, 61, 62]
@@ III.1 Conical space-time
where $\mu_{G}$ is the string tension. The topology of this space is nontrivial in the sense that it is locally flat everywhere except at the origin where the curvature has a singularity. This singularity is precisely the reason why null-geodesics are all deflected by the same angle,
@@ III.2 Full-wave solution
Although Eq. (8) is self-consistent and enables the computation of the transmission factor over the full range of parameters, it is helpful to examine certain limits—such as the geometrical optics (GO) and line-of-sight cases. These limits simplify the general expressions into more analytically tractable forms, making the physical interpretation more transparent.
@@ III.3 High frequency limit
At high frequencies, $ft_{\Delta}\gg 1$, the Fresnel integrals can be approximated using asymptotic expansion [63],
@@ III.5 Comparison with Point Mass Lens
Figure 4: Density plot of the transmission factor $|F|$ for a point-mass lens, shown as a function of the rescaled frequency $ft_{M}$ and source position $y=\theta/\theta_{\rm E}$. See Ref. [59] for details.
@@ IV Cosmic String Lensing Imprints and Detectability
We are interested in the behavior of the transmission factor within the frequency range accessible to the LVK network, which spans approximately from $f_{-}\approx 30\,$Hz to $f_{+}\approx 1\,$kHz [88, 89] The transmission factor for cosmic string lensing, given in Eq. (8), reaches its maximum amplification when the source is perfectly aligned with the line of sight ($y=0$) occurring at a characteristic frequency of approximately $f_{00}=(3/4)/t_{\Delta}$, as derived in Appendix B. At this frequency, the magnitude attains its peak value, $|F|_{\rm max}\approx 2.34$, representing the most prominent feature of wave-optics lensing [see Eq. (50)]. The amplification consists of a GO contribution factor of 2, corresponding to the two interfering rays, and an additional $\approx 0.34$ arising from the diffracted wave. For the lensing effect to be detectable by the LVK network, the condition $f_{00}<f_{+}$ must be satisfied. If $f_{00}$ significantly exceeds $f_{+}$, the amplification rapidly diminishes across all values of the parameter $y$ within the detector band, rendering the lensing signature increasingly difficult to detect. Figure 5 illustrates how the frequency at which the maximum amplification $|F|_{\mathrm{max}}$ occurs shifts to higher values as the string tension $\Delta$, or, equivalently, the lensing time delay $t_{\Delta}$, is decreased.
@@
which, for the LVK frequency range, simplifies to:
@@
To better understand the gravitational lensing effect, it is useful to visualize how it imprints on the gravitational waveform. The time-domain strain $h(t)$ of the lensed signal is obtained by taking the inverse Fourier transform of the frequency-domain waveform $\tilde{h}_{L}(f)$ modified by the transmission factor, as given in Eq. (24). The resulting lensed waveform is shown in Fig. 7. Correspondingly, Table 1 summarizes the values of the parameters, both for the BBH source and the lens, used to generate the figure. Here, we distinguish between two qualitatively different lensing regimes: (i) microlensing regime — when the time delay between the two images is small, the waveforms overlap and interfere, producing a characteristic beating pattern. (ii) strong lensing regime — when the time delay is large enough for the two lensed images to be well separated in time. In the microlensing case [Fig. 7(a)], the earliest signal consists of a beating pattern from the interference of both images, followed by the first merger peak. The second image arrives later without interference, keeping a shape identical to the original waveform. In the strong lensing case [Fig. 7(b)], both images are exact, time-shifted replicas of the original signal—a direct consequence of the topological nature of the lensing. The presence of two distinct merger peaks in both regimes is a unique signature of gravitational lensing, distinguishing it from other effects that can produce beating.
@@
Figure 7: Time-domain strain of a lensed GW (orange) compared to the unlensed waveform (blue). Two different lensing regimes are illustrated: (a) microlensing regime, where the two images overlap and interfere, producing a beating pattern. (b) strong lensing regime, with two images separated in time. The time delays are indicated by red lines and are reported in Table 1.
V Mismodelling biases and identification of CS signatures.
For the detection of compact binary coalescence signals, most GW search pipelines rely on matched filtering [91, 89]. However, the absence of lensed waveforms in the template banks used in matched filtering introduces a systematic bias. In this work, we investigate the impact of this bias in the case of a CS lens. We disentangle two sources of detection bias when analyzing lensed signals with unlensed template banks. First, we quantify the loss of matched-filter signal-to-noise ratio (SNR) arising from the waveform mismatch between the true lensed signal and the unlensed templates. Second, we assess the additional reduction in detection significance caused by the $\chi^{2}$ signal-consistency test [92], which is employed in the PyCBC pipeline [93, 94, 95] for glitch mitigation. Reference [84] explicitly accounts for this effect via full injection campaigns of PML–lensed signals analyzed with unlensed templates. Furthermore, we investigate the extent to which CS–lensed signals can be distinguished from those lensed by a PML and from unlensed signals by computing the odds ratio. To do this we estimate the prior ratio from detection rates and compute the evidence ratio from parameter estimations using the Bilby framework.
@@ V.1 Matched filtering and consistency tests
A matched filtering algorithm typically searches for a transient signal by computing the correlation between the detector data and a bank of template waveforms. This correlation is given by [96, 97],
@@
where $d(f)$ is the data employed in the search, $h_{j}(f)$ is the $j$-th template in the bank, and $S_{n}(f)$ is the detector’s noise power spectral density (PSD). The matched-filter output $z_{j}(t)$ is a complex time series, and its magnitude $\rho(t)=|z_{j}(t)|$, defines the SNR. A candidate event is identified when $\rho(t)$ exceeds a predefined threshold. Under the assumption of stationary Gaussian noise, the SNR is maximized when the template $h(f)$ matches the true signal [91].
@@
where $\varphi$ denotes the signal phase, $t_{c}$ is the coalescence time, and the subscript $j$ refers to the $j$-th waveform in the template bank. A mismatch of zero corresponds to perfect agreement between the signal and the template, yielding the optimal SNR, $\rho_{\rm opt}$, which serves as an upper bound:
@@
Signal consistency tests, such as the $\chi^{2}$ test, are traditionally used to suppress non-astrophysical transients, or glitches, that may mimic real signals [98]. However, their utility extends beyond glitch rejection. These tests also evaluate whether the signal’s contribution to the matched-filter SNR is consistent across frequency bands, helping to distinguish true signals from structured noise [92]. Additionally, they provide a measure of how well the data matches the expected waveform, which can be used to improve event ranking, assess signal fidelity, and support multi-detector coincidence analyses [99, 94].
@@
Figure 8: Impact of template mismatch and signal consistency tests on the SNR of cosmic string–lensed signals. Each row corresponds to one of the injections listed in Table 3, ordered by increasing injected SNR from top to bottom. The columns show, from left to right: (i) the mismatch between lensed and unlensed waveforms, (ii) the SNR loss due to the $\chi^{2}$ consistency test, quantified by the ratio $\hat{\rho}/\rho$, and (iii) the total SNR loss relative to the optimal injected value, given by $\hat{\rho}/\rho_{\rm opt}$.
V.2 Quantitative analysis of the lensing mismodelling bias
Using the PyCBC package, we inject three BBH signals generated with the IMRPhenomXPHM [90] waveform, produced with the default MSA spin prescription into simulated noise based on the Advanced LIGO PSD under the high-sensitivity settings of the T2000012-v2 technical document [^4] https://dcc.ligo.org/LIGO-T2000012/public. The parameters of our BBH waveforms are listed in Table 3. We then apply the CS transmission factor (8), exploring lensing parameters in the ranges $t_{\Delta}\in[10^{-3},1]\,\mathrm{ms}$ and $y\in[10^{-3},1]$, over a $30\times 30$ equally spaced grid. For each lensed signal, we compute the mismatch with the corresponding unlensed waveform, the matched-filter SNR time series, the $\chi^{2}$ test, and the reweighted SNR. The results are linearly interpolated to produce a 2D visualization across the full range of lensing parameters.
Table 2: BBH signal parameters of the three injections used in this study. $\mathcal{M}$ is the chirp mass, $q$ is the mass ratio, and $M_{\rm T}$ is the total mass of the binary. $\chi_{iz}$ denotes the $z$ component of the dimensionless spin parameter for the $i$-th black hole (the other components are negligible). $\varphi_{c}$ is the phase at coalescence, while $\alpha$ and $\delta$ are the right ascension and declination angles, respectively. The injection SNR refers to the SNR obtained for the unlensed injection.

### v3

# Probing cosmic strings via gravitational-wave lensing

@@ Abstract
We present a framework for detecting gravitational-wave signals lensed by cosmic strings (CSs), addressing a key gap in current searches. CSs, whose detection would provide a unique probe of high-energy physics and the early Universe, possess distinct topological and geometric features that require a dedicated search strategy. Our approach employs a full-wave transmission factor, expressed analytically via Fresnel integrals, which captures the characteristic diffraction and interference effects of the conical spacetime around a straight CS. We contrast CS lensing with the well-studied point mass lens (PML) model, highlighting their fundamental differences: CS lensing depends on cosmological distances, string tension $\Delta$, and wavelength $\lambda$, and produces two non-amplified images set by the global conical geometry. In contrast, PML lensing is governed by the distance-independent ratio $\sim M_{Lz}/\lambda$, where $M_{Lz}$ represents the redshifted mass of the lens, with image properties derived from the lens equation. For BBH mergers lensed by CSs, we show that the waveforms exhibit a characteristic beating pattern or time-separated exact replicas. We derive a detectability bound on the string tension and, using Bayesian model selection, demonstrate that CS lensing is distinguishable from both unlensed and PML-lensed signals across a wide region of parameter space.
@@ I Introduction
Cosmic strings (CSs) are one-dimensional topological defects predicted to arise during spontaneous symmetry-breaking phase transitions in the early Universe, within a wide range of quantum field theories, including grand unified theories and string theory [1, 2, 3, 4, 5]. Their detection would provide a unique observational window into high-energy physics and the early Universe, potentially validating fundamental theoretical models. Advanced LIGO [6], Advanced Virgo [7], and KAGRA [8] observatories form an international network of ground-based gravitational-wave (GW) detectors. Together, they constitute the LIGO-Virgo-KAGRA (LVK) Collaboration, which has recently undertaken its fourth observing run (O4) [9]. The continuous improvement of detector sensitivity and global coordination among these observatories significantly enhance the prospects of detecting a broad range of GW sources, including potential signals from CSs. The LVK Collaboration has been actively searching for GW signals emitted by CS loops and kinks [10, 11, 12]. When a network of CSs is assumed, their collective emission may contribute to the stochastic GW background [13, 14, 15, 16, 17]. In addition, several studies have explored the possibility of identifying individual GW bursts from CSs [18, 19].
@@
To address this gap, we introduce a comprehensive framework tailored for detecting gravitational lensing signatures from CSs. Central to our approach is a full-wave transmission factor that incorporates interference and diffraction effects from the conical spacetime geometry of CSs, providing a compact analytical expression in terms of Fresnel integrals readily available in scientific libraries such as Python’s scipy. This formulation enables the efficient construction of matched-filter templates and their straightforward implementation in data-analysis pipelines, laying the groundwork for targeted searches of CS-induced lensing effects in GW data. Focusing on the case of BBH mergers lensed by straight CSs, we characterize the resulting observational signatures, assess their detectability with current and future GW observatories, and examine potential mismodeling biases, matched-filtering strategies, and consistency tests to provide a quantitative evaluation of their impact on lensing detectability. Furthermore, we study the characteristic imprint left by CS lensing on detected GW and its discernibility from other, more common GW sources by computing the odds ratio PML lensed waveforms or unlensed waveforms.
This paper is organized as follows. In Sec. II, we review the main aspects of wave-optics effects in gravitational lensing, distinguishing between compact lenses and cosmic strings, and introduce the key parameters that characterize each regime. In Sec. III, we begin with the conical spacetime geometry and present the full-wave solution for the transmission factor, parametrized by two dimensionless quantities: the scaled GW frequency—set by the string tension—and the source offset relative to the line of sight. Unlike compact lenses, the thin-lens approximation does not apply to cosmic strings; no lens equation is required, as image positions are dictated directly by the global conical geometry and the string tension. Finally, we compare cosmic string lensing with the point mass lens, emphasizing the fundamentally different physical mechanisms and observational signatures. Section IV discusses the observational imprints of cosmic string lensing and their detectability. In Sec. V, we analyze potential selection biases, including matched filtering strategies and consistency checks, and provide a quantitative assessment of their impact on lensing detectability. Finally, Sec. VI summarizes our conclusions. Additional technical details and estimate of the detection rate are provided in the appendices.
@@ III The transmission factor
III.1 Conical spacetime
Suppose that the string (referred to as the “lens”) is positioned very close to the line of sight between the source of the GW and the detector. For simplicity, we assume that the distance $d_{LS}$ between the GW source and the string is much longer than $d_{L}$ – the distance between the string and the detector. In this case the incident wave is just a plane wave and the reduced distance $\mathcal{D}\approx d_{L}$. Additionally, the string is considered to be orthogonal to the line of sight. This assumption does not preclude the application of our analysis to the possibility of inclination.
Consider the spacetime metric for a static cylindrically symmetric cosmic string lying along the $z$ axis [61, 62, 63]
@@ III.1 Conical spacetime
where $\mu_{G}$ is the string tension. The topology of this space is nontrivial in the sense that it is locally flat everywhere except at the origin where the curvature has a singularity. This singularity is precisely the reason why null geodesics are all deflected by the same angle,
@@ III.2 Full-wave solution
Although Eq. (8) is self-consistent and enables the computation of the transmission factor over the full range of parameters, it is helpful to examine certain limits—such as the GO and line-of-sight cases. These limits simplify the general expressions into more analytically tractable forms, making the physical interpretation more transparent.
@@ III.3 High frequency limit
At high frequencies, $ft_{\Delta}\gg 1$, the Fresnel integrals can be approximated using the asymptotic expansion [64],
@@ III.5 Comparison with Point Mass Lens
Figure 4: Density plot of the transmission factor $|F|$ for a point mass lens, shown as a function of the rescaled frequency $ft_{M}$ and source position $y=\theta/\theta_{\rm E}$. See Ref. [60] for details.
@@ IV Cosmic String Lensing Imprints and Detectability
We are interested in the behavior of the transmission factor within the frequency range accessible to the LVK network, which spans approximately from $f_{-}\approx 30\,$Hz to $f_{+}\approx 1\,$kHz [89, 90] The transmission factor for cosmic string lensing, given in Eq. (8), reaches its maximum amplification when the source is perfectly aligned with the line of sight ($y=0$) occurring at a characteristic frequency of approximately $f_{00}=(3/4)/t_{\Delta}$, as derived in Appendix B. At this frequency, the magnitude attains its peak value, $|F|_{\rm max}\approx 2.34$, representing the most prominent feature of wave-optics lensing [see Eq. (52)]. The amplification consists of a GO contribution factor of 2, corresponding to the two interfering rays, and an additional $\approx 0.34$ arising from the diffracted wave. For the lensing effect to be detectable by the LVK network, the condition $f_{00}<f_{+}$ must be satisfied. If $f_{00}$ significantly exceeds $f_{+}$, then the amplification rapidly diminishes across all values of the parameter $y$ within the detector band, rendering the lensing signature increasingly difficult to detect. Figure 5 illustrates how the frequency at which the maximum amplification $|F|_{\mathrm{max}}$ occurs shifts to higher values as the string tension $\Delta$, or, equivalently, the lensing time delay $t_{\Delta}$, is decreased.
@@
which, for the LVK frequency range, simplifies to
@@
To better understand the gravitational lensing effect, it is useful to visualize how it imprints on the gravitational waveform. The time-domain strain $h(t)$ of the lensed signal is obtained by taking the inverse Fourier transform of the frequency-domain waveform $\tilde{h}_{L}(f)$ modified by the transmission factor, as given in Eq. (24). The resulting lensed waveform is shown in Fig. 7. Correspondingly, Table 1 summarizes the values of the parameters, both for the BBH source and the lens, used to generate the figure. Here, we distinguish between two qualitatively different lensing regimes: (i) microlensing regime—when the time delay between the two images is small, the waveforms overlap and interfere, producing a characteristic beating pattern, and (ii) strong lensing regime—when the time delay is large enough for the two lensed images to be well separated in time. In the microlensing case [Fig. 7(a)], the earliest signal consists of a beating pattern from the interference of both images, followed by the first merger peak. The second image arrives later without interference, keeping a shape identical to the original waveform. In the strong lensing case [Fig. 7(b)], both images are exact time-shifted replicas of the original signal—a direct consequence of the topological nature of the lensing. The presence of two distinct merger peaks in both regimes is a unique signature of gravitational lensing, distinguishing it from other effects that can produce beating.
@@
Figure 7: Time-domain strain of a lensed GW (orange) compared to the unlensed waveform (blue). Two different lensing regimes are illustrated: (a) microlensing regime, where the two images overlap and interfere, producing a beating pattern, and (b) strong lensing regime, with two images separated in time. The time delays are indicated by red lines and are reported in Table 1.
V Mismodeling biases and identification of CS signatures.
For the detection of compact binary coalescence signals, most GW search pipelines rely on matched filtering [92, 90]. However, the absence of lensed waveforms in the template banks used in matched filtering introduces a systematic bias. In this work, we investigate the impact of this bias in the case of a CS lens. We disentangle two sources of detection bias when analyzing lensed signals with unlensed template banks. First, we quantify the loss of matched-filter signal-to-noise ratio (SNR) arising from the waveform mismatch between the true lensed signal and the unlensed templates. Second, we assess the additional reduction in detection significance caused by the $\chi^{2}$ signal-consistency test [93], which is employed in the PyCBC pipeline [94, 95, 96] for glitch mitigation. Reference [85] explicitly accounts for this effect via full injection campaigns of PML-lensed signals analyzed with unlensed templates. Furthermore, we investigate the extent to which CS-lensed signals can be distinguished from those lensed by a PML and from unlensed signals by computing the odds ratio. To do this we estimate the prior ratio from detection rates and compute the evidence ratio from parameter estimations using the Bilby framework.
@@ V.1 Matched filtering and consistency tests
A matched filtering algorithm typically searches for a transient signal by computing the correlation between the detector data and a bank of template waveforms. This correlation is given by [97, 98]
@@
where $d(f)$ is the data employed in the search, $h_{j}(f)$ is the $j$th template in the bank, and $S_{n}(f)$ is the detector’s noise power spectral density (PSD). The matched-filter output $z_{j}(t)$ is a complex time series, and its magnitude $\rho(t)=|z_{j}(t)|$, defines the SNR. A candidate event is identified when $\rho(t)$ exceeds a predefined threshold. Under the assumption of stationary Gaussian noise, the SNR is maximized when the template $h(f)$ matches the true signal [92].
@@
where $\varphi$ denotes the signal phase, $t_{c}$ is the coalescence time, and the subscript $j$ refers to the $j$th waveform in the template bank. A mismatch of zero corresponds to perfect agreement between the signal and the template, yielding the optimal SNR, $\rho_{\rm opt}$, which serves as an upper bound:
@@
Signal consistency tests, such as the $\chi^{2}$ test, are traditionally used to suppress non-astrophysical transients, or glitches, that may mimic real signals [99]. However, their utility extends beyond glitch rejection. These tests also evaluate whether the signal’s contribution to the matched-filter SNR is consistent across frequency bands, helping to distinguish true signals from structured noise [93]. Additionally, they provide a measure of how well the data matches the expected waveform, which can be used to improve event ranking, assess signal fidelity, and support multidetector coincidence analyses [100, 95].
@@
[…]
V.2 Quantitative analysis of the lensing mismodeling bias
Using the PyCBC package, we inject three BBH signals generated with the IMRPhenomXPHM [91] waveform, produced with the default multi-scale analysis spin prescription into simulated noise based on the Advanced LIGO PSD under the high-sensitivity settings of the T2000012-v2 technical document [^4] https://dcc.ligo.org/LIGO-T2000012/public. The parameters of our BBH waveforms are listed in Table 2. We then apply the CS transmission factor (8), exploring lensing parameters in the ranges $t_{\Delta}\in[10^{-3},1]\,$s and $y\in[10^{-2},1]$, over a $30\times 30$ equally spaced grid. This range of parameters, together with the chosen signal duration, has been selected to ensure that the lensing effects remain in the microlensing regime [Fig.7(a)]. In the strong-lensing regime, the original waveform is not significantly modulated; instead, two time-delayed copies of the signal are produced [Fig.7(b)]. In this situation, the mismatch between the unlensed and lensed waveforms is null, so no mismodeling bias arises. For each lensed signal, we compute the mismatch with the corresponding unlensed waveform, the matched-filter SNR time series, the $\chi^{2}$ test, and the reweighted SNR. The results are linearly interpolated to produce a 2D visualization across the full range of lensing parameters.
Table 2: BBH signal parameters of the three injections used in this study. $\mathcal{M}$ is the chirp mass, $q$ is the mass ratio, and $M_{\rm T}$ is the total mass of the binary. $\chi_{iz}$ denotes the $z$ component of the dimensionless spin parameter for the $i$th black hole (the other components are negligible). $D_{L}$ is the luminosity distance to the source, $\varphi_{c}$ is the phase at coalescence, while $\alpha$ and $\delta$ are the right ascension and declination angles, respectively. The injection SNR refers to the SNR obtained for the unlensed injection.